will a precipitate occur if 250 ml of 1.5 x 10^.3 M NaCl is mixed with 250 mL of 2.0 x 10^.7 M AgNO3? Ksp of AgCl = 1.6 x 10^-10
1
Expert's answer
2013-03-29T12:25:09-0400
The equation for the chemical reaction is

NaCl(aq) + AgNO3(aq) = AgCl(s) + NaNO3(aq)

Ksp = [Ag+] x [Cl-] = 1.6 x 10^-10
From the equation for Ksp:
[Ag+] = [Cl-] = (1.6 x 10^-10) ^0.5 = 1.3 x 10^-5 M

The amount of NaCl is
n(NaCl) = C(NaCl) x V(L) = 1.5 x 10^-3 x 0.25 = 3.75 x 10^-4 mol

The amount of AgNO3 is
n(AgNO3) = C(AgNO3) x V(L) = 2.0 x 10^-7 x 0.25 = 5.0 x 10^-8 mol

The precipitate will occur if the product of [Ag+] and [Cl-] will be more or equal to Ksp:
[Ag+] = n(Ag+) / V(of new solution) = 5.0 x 10^-8 / (0.25 + 0.25) = 1.0 x 10^-7 M
[Cl-] = n(Cl-)/V(of new solution) = 3.75 x 10^-4 / (0.25 + 0.25) = 7.5 x 10^-4 M
[Ag+] x [Cl-] = 1.0 x 10^-7 x 7.5 x 10^-4 = 7.5 x 10^-11 it is less than Ksp, that's why the precipitate will not occur

Answer: The precipitate won't occur.